OpenStudy (anonymous):

how to solve 20x-7=5=16x

OpenStudy (anonymous):

was there a typo?

OpenStudy (anonymous):

20x - 7 = 5 - 16x 4x = 12 x = 3.

OpenStudy (anonymous):

No, wait that's wrong... 20x - 7 = 5 -16x 36x = 12 3x = 1. x = 1/3.